# Morse-Bott inequalities, Topology Change and Cobordisms to Nothing

### v1

@@ 1 Introduction
One of the most surprising aspects of superstring theories is the prediction of $D=10$ as the critical dimension ($D=26$ for the bosonic string) of spacetime. In order to match the observed $d=4$ of our universe, it is necessary to somehow get rid of the $n=D-d=6$ extra directions. A straightforward way to do this is by compactifying them, in such a way that they take values on a compact $n$-dimensional manifold $\mathcal{C}_{n}$.
@@
In the presence of such many vacua/low energy descriptions, the decay or transition into another is a general expectation, associated for example to different local minima in flux compactifications Kachru:2002gs ; Blanco-Pillado:2009lan , as well as topology changing transitions to a different compact manifold Atiyah:2000zz ; Acharya:2004qe , such as flops Aspinwall:1993yb ; Witten:1993yc or conifolds Strominger:1995cz ; Greene:1995hu . Generally speaking, topology changes are a generic expectation of Quantum Gravity, Giddings:1987cg ; Greene:1990ud ; Giveon:1994fu ; Witten:1998zw ; Garcia-Etxebarria:2018ajm ; Demulder:2023vlo .
@@
Another type of instability was shown by Witten Witten:1981gj , where the decay does not occur to a different vacuum, but rather to nothing, this is, to a configuration without spacetime. This construction, which will be reviewed in more detail in section 2.1, consists in a decay channel for a Minkowski vacuum with a single extra dimension compactified on a circle of radius $\mathsf{R}$. A sphere of size $\mathsf{R}$ pops up, with spacetime ending smoothly on its surface, in such a way that its interior is non-accessible, with “nothing” (not even spacetime) being inside. The bubble expands with a velocity asymptotic to that of light, and as we approach its surface both the lower-dimensional curvature and the (normalized) scalar controlling the size of the the circle blow up. This type of vacuum decay, dubbed Bubble of Nothing (BoN) has been generalized to other compactification manifolds and internal ingredients Fabinger:2000jd ; Dine:2004uw ; Horowitz:2007pr ; Yang:2009wz ; Blanco-Pillado:2010xww ; Blanco-Pillado:2010vdp ; Brown:2010mf ; Blanco-Pillado:2011fcm ; Brown:2011gt ; deAlwis:2013gka ; Brown:2014rka ; Blanco-Pillado:2016xvf ; Ooguri:2017njy ; Dibitetto:2020csn ; GarciaEtxebarria:2020xsr ; Bomans:2021ara ; Draper:2021ujg ; Draper:2021qtc ; Petri:2022yhy ; Bandos:2023yyo ; Ookouchi:2024tfz ; Heckman:2024zdo .
@@
Bubbles of nothing and other spacetime ending configurations, such as Hořava-Witten walls Horava:1995qa ; Horava:1996ma , might seem at first a curiosity without further implications. However, as shown in McNamara:2019rup , this is not the case at all. The Swampland program Vafa:2005ui ; Brennan:2017rbf ; Palti:2019pca ; vanBeest:2021lhn ; Grana:2021zvf ; Harlow:2022ich ; Agmon:2022thq , studies the set of constrains that an EFT with a consistent UV completion to Quantum Gravity must fulfill. One of the most established Swampland Conjectures is the No Global Symmetries conjecture Banks:1988yz ; Kallosh:1995hi ; Banks:2010zn , which postulates that there are no global charges in quantum gravity, so that any symmetry must be gauged or broken. A specific case of these global symmetries are topological ones. Since we expect some topology changes to be dynamically allowed in a theory of Quantum Gravity, this notion of topology change is better encapsulated by that of cobordism class. Two $k$-dimensional compact manifolds $\mathcal{M}_{k}$ and $\mathcal{N}_{k}$ are said to be cobordant if there exists a $(k+1)$ manifold $\mathcal{B}_{k+1}$ such that $\partial\mathcal{B}_{k+1}=\mathcal{M}\sqcup\overline{\mathcal{N}}$. $\mathcal{M}_{k}$ and $\mathcal{N}_{k}$ are then said to be (co)bordant and $\mathcal{B}_{k+1}$ is said to be a bordism between them. From a physical point, if compactifying our higher-dimensional theory on $\mathcal{M}_{k}$ and $\mathcal{N}_{k}$ results two distinct lower-energy descriptions, EFT1 and EFT2, then traversing $\mathcal{B}_{k+1}$ can be thought as crossing the domain wall relating the two. This is illustrated in Figure 1.
@@
Figure 1: Sketch of different bordisms between $n$-manifolds. In subfigure 1(a) the bordism $\mathcal{B}_{n+1}$ has as components of its boundary $\mathcal{M}_{n}$ and $\mathcal{N}_{n}$, and can be interpreted as a domain wall between the two EFTs resulting from compactification on each of the manifolds. In subfigure 1(b) $\mathcal{M}_{n}=\partial\mathcal{B}_{n+1}$, the bordism connects the EFT to nothing, serving as a boundary of spacetime, which from the lower-dimensional perspective we will call End of the World brane in section 2.2.
@@
Two compactifications that can be interpolated through a bordism or, in other words, for which there is a dynamically allowed topology change between them, are part of the same cobordism class. The topological charge associated to said classes is global rather than gauged McNamara:2019rup , as one can take some spacetime $\mathcal{X}_{d}$, remove some submanifold $\mathcal{M}_{k}\subset\mathcal{X}_{d}$, and glue along it some other $\mathcal{N}_{k}$ in the same cobordism class as $\mathcal{M}_{k}$. The resulting spacetime is cobordant to the original one, but far from the region where we have changed the local topology, both look the same, and the topology of this defect introduced cannot be inferred. Per the No Global Symmetries conjecture, this cannot occur in a gravitational theory with a consistent UV completion. The solution to this problem is this topological charge being trivial, or in other words, by requiring that our compactification manifold must be the boundary of an appropriate bordism, $\mathcal{M}_{n}=\partial\mathcal{B}_{n+1}$. This is known as the Swampland Cobordism Conjecture McNamara:2019rup .
@@
While the above might seem like a pretty abstract discussion about compactifications of QG, it has a radical implication: since then all compactifications of QG, say string theory, must be in the same cobordism class, they should be connected through some domain wall to nothing, this is, have spacetime ending boundary such as Witten’s BoN. [^1] As discussed in section 2.1, an extra requirement will be that such decay is dynamically allowed, this being a general expectation in non-SUSY compactifications. From the bordism point of view, this corresponds with the internal manifold smoothly capping off to a point, after which spacetime ends. If more stringy, singular objects are allowed, such as O-planes, then this is not necessarily the case, as the compact manifold itself can have boundaries, which themselves serve as domain walls to nothing. More exotic defects might be needed to kill cobordism classes to the trivial one, which has result in the prediction of new, often non-supersymmetric objects McNamara:2019rup ; Montero:2020icj ; Dierigl:2022reg , but their inclusion in the bordism is likely to result in singular solutions.
@@
Controlled solutions, solving the equations of motion, and describing explicit bordisms to nothing are rare in the literature, and usually correspond with simple compactification manifolds such as spheres or tori Fabinger:2000jd ; Dine:2004uw ; Horowitz:2007pr ; Yang:2009wz ; Blanco-Pillado:2010xww ; Blanco-Pillado:2010vdp ; Brown:2010mf ; Blanco-Pillado:2011fcm ; Brown:2011gt ; deAlwis:2013gka ; Brown:2014rka ; Blanco-Pillado:2016xvf ; Ooguri:2017njy ; Dibitetto:2020csn ; Bomans:2021ara ; Draper:2021ujg ; Draper:2021qtc ; Petri:2022yhy ; Bandos:2023yyo ; Ookouchi:2024tfz ; Heckman:2024zdo . The few explicit examples for more complicated cases without the need of stringy defects are quite involved GarciaEtxebarria:2020xsr . Smooth solutions are crucial from the phenomenological point of view, as they can be approached from the supergravity approximation. Furthermore, their nucleation rate (per unit volume) as a decay channel for the initial configuration will be dominant, and can be computed semiclassically $\Gamma_{\rm BoN}/{\rm Vol}_{d}\sim e^{\Delta S^{\rm(E)}}$ Coleman:1977py , with $\Delta S^{\rm(E)}$ the difference in Euclidean action between the initial vacua and the BoN solution. Singular bordisms will have divergent terms in their Euclidean action, and thus will be suppressed.
@@
The goal of our paper is to take a first step in the construction of smooth bordisms into nothing, as well as those which might have singular terms, but for which nonetheless $\mathcal{B}_{n+1}$ is homeomorphic to a smooth manifold. We will call said bordisms smoothable. We will obtain bounds in the topology of $\mathcal{B}_{n+1}$ from that of $\mathcal{C}_{n}=\partial\mathcal{B}_{n+1}$, and understand how this changes as we move along the bordism. Precisely for compact manifolds $\mathcal{C}_{n}$ with an involved internal topology, featuring different internal cycles of various dimensions, these smooth bordisms to nothing $\mathcal{B}_{n+1}$ cannot be the complete manifold $\mathcal{C}_{n}$ as a whole smoothly shrinking to a point [^2] One could simply consider the cone $C(\mathcal{C}_{n})=(\mathcal{C}_{n}\times[0,1])/\sim$, with $(x,1)\sim(y,1)$ for all $x,\,y\in\mathcal{C}_{n}$, but it can be shown that $C(\mathcal{C}_{n})$ is not a topological manifold unless $H_{k}(\mathcal{C}_{n},\mathbb{Z})\simeq\mathbb{Z}$ for $k=0,\,n$ and 0 otherwise, i.e., $\mathcal{C}_{n}$ is a homology sphere., but rather has a more involved structure, with the different internal cycles shrinking before the final tip into nothing. This corresponds with intermediate topology changes, which will be associated with a series of domain walls before, or “coating”, the final boundary to nothing. Our approach will be able to give us information about the internal structure of both Bubbles of Nothing that dynamically nucleate and, as well as more general fixed End of the World branes in interpolating from our EFT theory into nothing.
@@
As it is easy to imagine, explicit solutions featuring these intermediate topology changes are quite involved. We will not attempt to write those, but simply obtain qualitative information about the structure of smooth(able) bordisms into nothing. The tool for this will be Morse theory, interpreting the bordism direction in $\mathcal{B}_{n+1}$ as our Morse function, whose critical points we will understand as the loci where the internal topology changes. As a matter of fact, Morse theory is nothing new in high energy theoretical physics, with the seminal work by Witten Witten:1982im giving a supersymmetric system interpretation, which allowed him to derive a version of the so-called Morse inequalities we will use in this work, relating the set of critical points of a function with the topology of the manifold it is defined over. [^3] Morse theory had already been considered in the literature Sorkin:1989ea ; Louko:1995jw ; Dowker:1997kc ; Dowker:1997hj ; Dowker:1999wu ; Garcia-Heveling:2022fkf for spacetime topology changes. Unlike in our case, in which these changes occur as we move spatially towards the end of our spacetime (be it the tip of the BoN or more generally the End of the World brane, see section 2), there the topology changes occur in the time direction, with the Lorentzian metric vanishing precisely at the critical points of the Morse function with which it is constructed. The similarities between our approach and theirs end here, as we will not need to worry about problems of causality, the main focus of attention of said references. Our work goes in line with recent approaches trying to infer local information of the EFT from topological global properties Lust:2024aeg . We will also be able to use the Morse theory tools to obtain qualitative properties of configurations where more than one bordism to nothing is present.
@@ 2.1 Bubbles of Nothing
The simplest (and coincidentally first found to appear in the literature) dynamical bordism into nothing is Witten’s Bubble of Nothing Witten:1981gj , presented here for an arbitrary number of spacetime (non-compact) dimensions. To study this non-perturbative instability of Kaluza-Klein vacuum, consider a circle compactification to a $d$-dimensional Minkowski, with global spacetime $X_{d+1}=\mathbb{R}^{1,d}\times\mathbb{S}^{1}$, with $\mathbb{S}^{1}$ of radius $\mathsf{R}$. This instability can be constructed from a Schwarzschild instanton $\hat{\mathcal{S}}_{d+1}$ in the Euclidean continuation of our spacetime, $X_{d+1}^{\rm(E)}=\mathbb{R}^{d+1}\times\mathbb{S}^{1}$, which will have metric
$\mathrm{d}s_{d+1}^{2}=r^{2}\mathrm{d}\Omega_{d-1}^{2}+\frac{\mathrm{d}r^{2}}{1-\left(\frac{R}{r}\right)^{d-2}}+\mathsf{R^{2}}\left[1-\left(\frac{R}{r}\right)^{d-2}\right]\mathrm{d}\theta^{2}\;,\quad\text{with }\theta\in[0,2\pi)\,,$ (2.1)
@@
However, precisely the fascinating observation made in Witten:1981gj is that the new time and radial coordinates $t$ and $x$ do not cover the full original $X_{d+1}$, only those parts with $x^{2}-t^{2}=r^{2}\geq R^{2}$, with $r=R$ corresponding to the wall of the bubble at the time of nucleation $t=0$. The “inside” of the $x^{2}-t^{2}=R^{2}$ hyperboloid is now removed from spacetime. The size of the extra dimension
$\hat{\mathsf{R}}(r)=\mathsf{R}\sqrt{1-\left(\frac{R}{r}\right)^{d-2}}$ (2.4)
@@
Figure 2: Euclidean sketch of the $d$-dimensional version of Witten’s BoN (together with the Wick rotated KK Minkowski vacuum), with the macroscopic dimensions being the horizontal gray direction depicted in gray. Notice that far from the BoN surface, both spaces are locally identical.
@@
As shown in Figure 2, the Euclidean solution has global topology $\hat{\mathcal{S}}_{d+1}\simeq\mathbb{S}^{d-1}\times_{\rm W}\mathbb{D}^{2}\simeq\mathbb{S}^{d-1}\times\mathbb{R}^{2}$, where ×$\times_{\rm W}$ is a warped product and $\mathbb{D}^{n}=\{x\in\mathbb{R}^{n}\,:\,\|x\|\leq 1\}$ is the $n$-disc, with indeed $\partial\mathbb{D}^{2}=\mathbb{S}^{1}$.
@@
Already in the original appearance of Witten’s BoN Witten:1981gj a possible topological obstruction to the KK vacuum decay was postulated in the presence of fermions. Being $\hat{\mathcal{S}}_{d+1}$ simply connected, there is a unique spinor structure, which here it is set by the asymptotics far from the bubble wall, where space locally looks like $\mathbb{R}^{d}\times\mathbb{S}^{1}$. Under a $2\pi$ rotation along $\mathbb{S}^{1}$, spinors can have either symmetric or antisymmetric boundary conditions. However, if we want to be able to extend the solution to the complete $\hat{\mathcal{S}}_{d-1}$, the spin structure must be extended to the whole $\mathbb{D}^{2}$. Since this is contractible, only antiperiodic spin structure makes sense for the BoN solution to exist. Since this accounts for non-covariantly constant spinors, SUSY must be broken either in the higher-dimensional theory or explicitly through the imposition of Scherk-Schwarz boundary conditions Scherk:1978ta ; Scherk:1979zr ; Kounnas:1989dk when compactifying.
@@
As for whether the vacuum decay is energetically allowed, the ADM mass computation shows (2.2) to have 0 energy (same as $X_{d+1}=\mathbb{R}^{1,d-1}\times\mathbb{S}^{1}$). The Positive Energy Theorem Schon:1979rg ; Witten:1981mf , i.e., that every non-flat solution of Einstein’s equations asymptoting to Minkowski has positive energy (thus semiclassically preventing Minkowski decay) does not apply here when antiperiodic boundary conditions are used in $\mathbb{S}^{1}$, as the proof of the theorem relays on the existence of covariantly constant spinors. The decay rate (per unit volume per unit time) can be computed to be $\Gamma\sim\exp\left\{-2\pi^{d/2}\Gamma(\frac{d}{2})^{-1}(\mathsf{R}M_{{\rm Pl},d})^{d-2}\right\}$, with all contributions coming from the Gibbons–Hawking–York term (the BoN solution being Ricci-flat) and having $\Gamma\to 0$ for large radius.
@@
As mentioned above, the topology of the Euclidean instanton $\hat{\mathcal{S}}_{d+1}$ (2.1) is not $\mathbb{R}^{d+1}\times\mathbb{S}^{1}$, but rather $\mathbb{S}^{d-1}\times_{\rm W}\mathbb{D}^{2}$. As explained in GarciaEtxebarria:2020xsr , this can be made more explicit by rewriting the Euclidean solution 2.1 as
@@
$\mathrm{d}s^{2}_{d+1}=\mathcal{W}(\rho)^{2}\mathrm{d}\Omega_{d-1}^{2}+\mathrm{d}\rho^{2}+\hat{\mathsf{R}}(\rho)^{2}\mathrm{d}\theta^{2}\;,\quad\text{with }\rho\in[0,\infty)\,,\;\theta\in[0,2\pi)\;,$ (2.5)
@@
where $\mathcal{W}(\rho)$ is some warping function.444 The radial directions are related by $r=\mathsf{R}W(\rho)$, with the bubble located at $\rho=0$ and the warping fulfilling $W^{\prime}(\rho)=\mathsf{R}^{-1}\sqrt{1-W(\rho)^{-(d-2)}}\,,\quad W(0)=1\;\Longleftrightarrow{}_{2}F_{1}\left(\frac{1}{2},\frac{1}{2-d},\frac{3-d}{2-d}W(\rho)^{2-d}\right)W(\rho)=\frac{\rho}{\mathsf{R}}\;,$ (2.6) which for large $\rho$ asymptotics to $W(\rho)\approx\frac{\rho}{\mathsf{R}}$ and $r\approx\rho$, as expected. Precisely for constant angle $\mathbb{S}^{d-1}$ direction in (2.5) the metric simplifies to

### v2

@@ 1 Introduction
One of the most surprising aspects of superstring theories is the prediction of $D=10$ as the critical dimension ($D=26$ for the bosonic string) of spacetime. In order to match the observed $d=4$ of our universe, it is necessary to somehow get rid of the n=Dd=6n=D-d=6 extra directions. A straightforward way to do this is by compactifying them, in such a way that they take values on a compact $n$-dimensional manifold $\mathcal{C}_{n}$.
@@
In the presence of such many vacua/low energy descriptions, the decay or transition into another is a general expectation, associated for example to different local minima in flux compactifications Kachru et al. (2002); Blanco-Pillado et al. (2009), as well as topology changing transitions to a different compact manifold Atiyah et al. (2001); Acharya and Gukov (2004), such as flops Aspinwall et al. (1993); Witten (1993); Greene et al. (2001) or conifolds Strominger (1995); Greene et al. (1995). Generally speaking, topology changes are a generic expectation of Quantum Gravity, Giddings and Strominger (1988); Greene and Plesser (1990); Giveon et al. (1994); Witten (1998); García-Etxebarria and Montero (2019); Demulder et al. (2024).
@@
Another type of instability was shown by Witten Witten (1982a), where the decay does not occur to a different vacuum, but rather to nothing, this is, to a configuration without spacetime. This construction, which will be reviewed in more detail in section 2.1, consists in a decay channel for a Minkowski vacuum with a single extra dimension compactified on a circle of radius $\mathsf{R}$. A sphere of size $\mathsf{R}$ pops up, with spacetime ending smoothly on its surface, in such a way that its interior is non-accessible, with “nothing” (not even spacetime) being inside. The bubble expands with a velocity asymptotic to that of light, and as we approach its surface both the lower-dimensional curvature and the (normalized) scalar controlling the size of the circle blow up. This type of vacuum decay, dubbed Bubble of Nothing (BoN) has been generalized to other compactification manifolds and internal ingredients Fabinger and Horava (2000); Dine et al. (2004); Horowitz et al. (2008); Yang (2010); Blanco-Pillado and Shlaer (2010); Blanco-Pillado et al. (2010); Brown and Dahlen (2011); Blanco-Pillado et al. (2012); Brown and Dahlen (2012); de Alwis et al. (2013); Brown (2014); Blanco-Pillado et al. (2016); Ooguri and Spodyneiko (2017); Dibitetto et al. (2020); García Etxebarria et al. (2020); Bomans et al. (2022); Draper et al. (2021b, a); Petri (2022); Bandos et al. (2023); Ookouchi et al. (2024); Heckman and Hübner (2024).
@@
Bubbles of nothing and other spacetime ending configurations, such as Hořava-Witten walls Horava and Witten (1996b, a), might seem at first a curiosity without further implications. However, as shown in McNamara and Vafa (2019), this is not the case at all. The Swampland program Vafa (2005); Brennan et al. (2017); Palti (2019); van Beest et al. (2022); Graña and Herráez (2021); Harlow et al. (2023); Agmon et al. (2022), studies the set of constrains that an EFT with a consistent UV completion to Quantum Gravity must fulfill. One of the most established Swampland Conjectures is the No Global Symmetries conjecture Banks and Dixon (1988); Kallosh et al. (1995); Banks and Seiberg (2011), which postulates that there are no global charges in quantum gravity, so that any symmetry must be gauged or broken. A specific case of these global symmetries are topological ones. Since we expect some topology changes to be dynamically allowed in a theory of Quantum Gravity, this notion of topology change is better encapsulated by that of cobordism class. Two $k$-dimensional compact manifolds $\mathcal{M}_{k}$ and $\mathcal{N}_{k}$ are said to be cobordant if there exists a $(k+1)$ manifold $\mathcal{B}_{k+1}$ such that $\partial\mathcal{B}_{k+1}=\mathcal{M}\sqcup\overline{\mathcal{N}}$. $\mathcal{M}_{k}$ and $\mathcal{N}_{k}$ are then said to be (co)bordant and $\mathcal{B}_{k+1}$ is said to be a bordism between them. From a physical point, if compactifying our higher-dimensional theory on $\mathcal{M}_{k}$ and $\mathcal{N}_{k}$ results two distinct lower-energy descriptions, EFT1 and EFT2, then traversing $\mathcal{B}_{k+1}$ can be thought as crossing the domain wall relating the two. This is illustrated in Figure 1.
@@
Figure 1: Sketch of different bordisms between $n$-manifolds. In subfigure 1(a) the bordism $\mathcal{B}_{n+1}$ has as components of its boundary $\mathcal{M}_{n}$ and $\mathcal{N}_{n}$, and can be interpreted as a domain wall between the two EFTs resulting from compactification on each of the manifolds. In subfigure 1(b) $\mathcal{M}_{n}=\partial\mathcal{B}_{n+1}$, the bordism connects the EFT to nothing, serving as a boundary of spacetime, which from the lower-dimensional perspective we will call End of the World brane in section $2.$
@@
Two compactifications that can be interpolated through a bordism or, in other words, for which there is a dynamically allowed topology change between them, are part of the same cobordism class. The topological charge associated to said classes is global rather than gauged McNamara and Vafa (2019), as one can take some spacetime $\mathcal{X}_{d}$, remove some submanifold $\mathcal{M}_{k}\subset\mathcal{X}_{d}$, and glue along it some other $\mathcal{N}_{k}$ in a different cobordism class then that of $\mathcal{M}_{k}$. The resulting spacetime is not cobordant to the original one, but far from the region where we have changed the local topology, both look the same, and the topology of this defect introduced cannot be inferred. Per the No Global Symmetries conjecture, this cannot occur in a gravitational theory with a consistent UV completion. The solution to this problem is this topological charge being trivial, or in other words, by requiring that our compactification manifold must be the boundary of an appropriate bordism, $\mathcal{M}_{n}=\partial\mathcal{B}_{n+1}$. This is known as the Swampland Cobordism Conjecture McNamara and Vafa (2019).
@@
While the above might seem like a pretty abstract discussion about compactifications of QG, it has a radical implication: since then all compactifications of QG, say string theory, must be in the same cobordism class, they should be connected through some domain wall to nothing, this is, have spacetime ending boundary such as Witten’s BoN. [^1] As discussed in section 2.1, an extra requirement will be that such decay is dynamically allowed, this being a general expectation in non-SUSY compactifications. From the bordism point of view, this corresponds to the internal manifold smoothly capping off to a point, after which spacetime ends. If more stringy, singular objects are allowed, such as O-planes, then this is not necessarily the case, as the compact manifold itself can have boundaries, which themselves serve as domain walls to nothing. More exotic defects might be needed to kill cobordism classes to the trivial one, which has result in the prediction of new, often non-supersymmetric objects McNamara and Vafa (2019); Montero and Vafa (2021); Dierigl et al. (2023), but their inclusion in the bordism is likely to result in singular solutions.
@@
Controlled solutions, solving the equations of motion, and describing explicit bordisms to nothing are rare in the literature, and usually correspond to simple compactification manifolds such as spheres or tori Fabinger and Horava (2000); Dine et al. (2004); Horowitz et al. (2008); Yang (2010); Blanco-Pillado and Shlaer (2010); Blanco-Pillado et al. (2010); Brown and Dahlen (2011); Blanco-Pillado et al. (2012); Brown and Dahlen (2012); de Alwis et al. (2013); Brown (2014); Blanco-Pillado et al. (2016); Ooguri and Spodyneiko (2017); Dibitetto et al. (2020); Bomans et al. (2022); Draper et al. (2021b, a); Petri (2022); Bandos et al. (2023); Ookouchi et al. (2024); Heckman and Hübner (2024). The few explicit examples for more complicated cases without the need of stringy defects are quite involved García Etxebarria et al. (2020). Smooth solutions are crucial from the phenomenological point of view, as they can be approached from the supergravity approximation. Furthermore, their nucleation rate (per unit volume) as a decay channel for the initial configuration will be dominant, and can be computed semiclassically $\Gamma_{\rm BoN}/{\rm Vol}_{d}\sim e^{\Delta S^{\rm(E)}}$ Coleman (1977), with $\Delta S^{\rm(E)}$ the difference in Euclidean action between the initial vacua and the BoN solution. Singular bordisms will have divergent terms in their Euclidean action, and thus will be suppressed.
@@
The goal of our paper is to take a first step in the construction of smooth bordisms into nothing, as well as those which might have singular terms, but for which nonetheless $\mathcal{B}_{n+1}$ is homeomorphic to a smooth manifold. We will call said bordisms smoothable. We will obtain bounds in the topology of $\mathcal{B}_{n+1}$ from that of $\mathcal{C}_{n}=\partial\mathcal{B}_{n+1}$, and understand how this changes as we move along the bordism. Precisely for compact manifolds $\mathcal{C}_{n}$ with an involved internal topology, featuring different internal cycles of various dimensions, these smooth bordisms to nothing $\mathcal{B}_{n+1}$ cannot be the complete manifold $\mathcal{C}_{n}$ as a whole smoothly shrinking to a point. [^2] One could simply consider the cone $C(\mathcal{C}_{n})=(\mathcal{C}_{n}\times[0,1])/\sim$, with $(x,1)\sim(y,1)$ for all $x,\,y\in\mathcal{C}_{n}$, but said cone $C(\mathcal{C}_{n})$ is not a topological manifold unless $H_{k}(\mathcal{C}_{n},\mathbb{Z})\simeq\mathbb{Z}$ for k=0,nk=0,\,n and 0 otherwise, i.e., $\mathcal{C}_{n}$ is a homology sphere. Rather, they will have a more involved structure, with the different internal cycles shrinking before the final tip into nothing. This corresponds to intermediate topology changes, which will be associated with a series of domain walls before, or “coating”, the final boundary to nothing. Our approach will be able to give us information about the internal structure of both Bubbles of Nothing that dynamically nucleate and, as well as more general End of the World branes when interpolating from our EFT theory into nothing.
@@
As it is easy to imagine, explicit solutions featuring these intermediate topology changes are quite involved. We will not attempt to write those, but simply obtain qualitative information about the structure of smooth(able) bordisms into nothing. The tool for this will be Morse theory, interpreting the bordism direction in $\mathcal{B}_{n+1}$ as our Morse function, whose critical points we will understand as the loci where the internal topology changes. As a matter of fact, Morse theory is nothing new in high energy theoretical physics, with the seminal work by Witten Witten (1982b) giving a supersymmetric system interpretation to it, which allowed him to derive a version of the so-called Morse inequalities we will use in this work, relating the set of critical points of a function with the topology of the manifold it is defined over. [^3] Morse theory had already been considered in the literature Sorkin (1989); Louko and Sorkin (1997); Dowker and Garcia (1998); Dowker and Surya (1998); Dowker et al. (2000); García-Heveling (2023) for spacetime topology changes. Unlike in our case, in which these changes occur as we move spatially towards the end of our spacetime (be it the tip of the BoN or more generally the End of the World brane, see section 2), there the topology changes occur in the time direction, with the Lorentzian metric vanishing precisely at the critical points of the Morse function with which it is constructed. The similarities between our approach and theirs end here, as we will not need to worry about problems of causality, the main focus of attention of said references. Our work goes in line with recent approaches trying to infer local information of the EFT from topological global properties Lüst (2024). We will also be able to use the Morse theory tools to obtain qualitative properties of configurations where more than one bordism to nothing is present.
@@ 2.1 Bubbles of Nothing
The simplest (and coincidentally first found to appear in the literature) dynamical bordism into nothing is Witten’s Bubble of Nothing Witten (1982a), presented here for an arbitrary number of spacetime (non-compact) dimensions. To study this non-perturbative instability of Kaluza-Klein vacuum, consider a circle compactification to a $d$-dimensional Minkowski, with global spacetime $X_{d+1}=\mathbb{R}^{1,d}\times\mathbb{S}^{1}$, with $\mathbb{S}^{1}$ of radius $\mathsf{R}$. This instability can be constructed from a Schwarzschild instanton $\hat{\mathcal{S}}_{d+1}$ in the Euclidean continuation of our spacetime, $X_{d+1}^{\rm(E)}=\mathbb{R}^{d+1}\times\mathbb{S}^{1}$, which will have metric
$\mathrm{d}s_{d+1}^{2}=r^{2}\mathrm{d}\Omega_{d-1}^{2}+\frac{\mathrm{d}r^{2}}{1-\left(\frac{R}{r}\right)^{d-2}}+\mathsf{R}^{2}\left[1-\left(\frac{R}{r}\right)^{d-2}\right]\mathrm{d}\theta^{2}\;,\quad\text{with }\left\{\begin{array}[]{l}\theta\in[0,2\pi)\\ r\geq R\end{array}\right$.\,, (2.1)
@@
However, precisely the fascinating observation made in Witten (1982a) is that the new time and radial coordinates $t$ and $x$ do not cover the full original $X_{d+1}$, but only those parts with $x^{2}-t^{2}=r^{2}\geq R^{2}$, with $r=R$ corresponding to the wall of the bubble at the time of nucleation $t=0$. The “inside” of the $x^{2}-t^{2}=R^{2}$ hyperboloid is now removed from spacetime. The size of the extra dimension
$\hat{\mathsf{R}}(r)=\mathsf{R}\sqrt{1-\left(\frac{R}{r}\right)^{d-2}}\;,$ (2.4)
@@
Figure 2: Euclidean sketch of the $d$-dimensional version of Witten’s BoN (together with the Wick rotated $K$ Minkowski vacuum), with the macroscopic dimensions being the horizontal gray direction depicted in gray. Notice that far from the BoN surface, both spaces are locally identical.
@@
As shown in Figure 2, the Euclidean solution has global topology $\hat{\mathcal{S}}_{d+1}\simeq\mathbb{S}^{d-1}\times_{\rm W}\mathbb{D}^{2}\simeq\mathbb{S}^{d-1}\times\mathbb{R}^{2}$, where $\times_{\rm W}$ is a warped product and $\mathbb{D}^{n}=\{x\in\mathbb{R}^{n}\,:\,\|x\|\leq 1\}$ is the $n$-disc, with indeed $\partial\mathbb{D}^{2}=\mathbb{S}^{1}$.
@@
Already in the original appearance of Witten’s BoN Witten (1982a) a possible topological obstruction to the KK vacuum decay was postulated in the presence of fermions. Being $\hat{\mathcal{S}}_{d+1}$ simply connected, there is a unique spinor structure, which here it is set by the asymptotics far from the bubble wall, where space locally looks like $\mathbb{R}^{d}\times\mathbb{S}^{1}$. Under a $2\pi$ rotation along $\mathbb{S}^{1}$, spinors can have either symmetric or antisymmetric boundary conditions. However, if we want to be able to extend the solution to the complete $\hat{\mathcal{S}}_{d-1}$, the spin structure must be extended to the whole $\mathbb{D}^{2}$. Since this is contractible, only the antiperiodic spin structure makes sense for the BoN solution to exist. Since this accounts for non-covariantly constant spinors, SUSY must be broken either in the higher-dimensional theory or explicitly through the imposition of Scherk-Schwarz boundary conditions Scherk and Schwarz (1979b, a); Kounnas and Rostand (1990) when compactifying.
@@
As for whether the vacuum decay is energetically allowed, the ADM mass computation shows (2.2) to have 0 energy (same as $X_{d+1}=\mathbb{R}^{1,d-1}\times\mathbb{S}^{1}$). The Positive Energy Theorem Schon and Yau (1979); Witten (1981), i.e., that every non-flat solution of Einstein’s equations asymptoting to Minkowski has positive energy (thus semiclassically preventing Minkowski decay), does not apply here when antiperiodic boundary conditions are used in $\mathbb{S}^{1}$, as the proof of the theorem relays on the existence of covariantly constant spinors. The decay rate (per unit volume per unit time) can be computed to be $\Gamma\sim\exp\left\{-2\pi^{d/2}\Gamma(\frac{d}{2})^{-1}(\mathsf{R}M_{{\rm Pl},d})^{d-2}\right\}$, with all contributions coming from the Gibbons–Hawking–York term (the BoN solution being Ricci-flat) and having $\Gamma\to 0$ for large radius.
@@
As mentioned above, the topology of the Euclidean instanton $\hat{\mathcal{S}}_{d+1}$ (2.1) is not $\mathbb{R}^{d+1}\times\mathbb{S}^{1}$, but rather $\mathbb{S}^{d-1}\times_{\rm W}\mathbb{D}^{2}$. As explained in García Etxebarria et al. (2020), this can be made more explicit by rewriting the Euclidean solution 2.1 as
@@
$\mathrm{d}s^{2}_{d+1}=W(\rho)^{2}\mathsf{R}^{2}\mathrm{d}\Omega_{d-1}^{2}+\mathrm{d}\rho^{2}+\hat{\mathsf{R}}(\rho)^{2}\mathrm{d}\theta^{2}\;,\quad\text{with }\rho\in[0,\infty)\,,\;\theta\in[0,2\pi)\;,$ (2.5)
@@
where $W(\rho)$ is some warping function. [^4] The radial directions are related by $r=\mathsf{R}W(\rho)$, with the bubble located at $\rho=0$ and the warping fulfilling $W^{\prime}(\rho)=\mathsf{R}^{-1}\sqrt{1-W(\rho)^{-(d-2)}}\,,\quad W(0)=1\;\Longleftrightarrow{}_{2}F_{1}\left(\frac{1}{2},\frac{1}{2-d},\frac{3-d}{2-d}W(\rho)^{2-d}\right)W(\rho)=\frac{\rho}{\mathsf{R}}\;,$ (2.6) which for large $\rho$ asymptotics to $W(\rho)\approx\frac{\rho}{\mathsf{R}}$ and $r\approx\rho$, as expected. Precisely for constant angle $\mathbb{S}^{d-1}$ direction in (2.5) the metric simplifies to